Wednesday, June 18, 2014

Night Log 2014-06-17

SA: Brent
SO: Thea
Others: Encarni, Eben, Luke, Eric, Christopher


Heavy cloud, tracker problem debugging (now resolved) and high humidity limited observing possibilities tonight.

Data taken for:

2014-1-RSA-001 P0 RSS [PSRB1259-63 during the 2014 periastron passage]
2014-1-DC-002 P3 RSS [Finding the redshift and ionization of WISE-selected QSOs]
2014-1-RSA_OTH-005 P4 RSS [Spectral study of the evolution of lenticular galaxies.]
2014-1-RSA_OTH_UKSC-001 P3 RSS [Long term monitoring of circumstellar disk in Be X-ray binaries]

No comments:

Post a Comment